OpenStudy (anonymous):

Calculate the average rate of change for the function f(x)=-x^4+4x^3-2x^2+x+1, from x=0 to x=1. 0 1 2 7

OpenStudy (harsha19111999):

Substitute x = 0 and x=1 in the given equation and see

OpenStudy (anonymous):

So 0=o^4+4(0)^3-2(0)^2+0+1 1=1^4+4(1)^3-2(1)^2+1+1

OpenStudy (harsha19111999):

for f(1)=1 , for f(1)=3

geerky42 (geerky42):

\(\dfrac{f(1)-f(0)}{1-0}\)

OpenStudy (anonymous):

So the answer is 1?

OpenStudy (harsha19111999):

Hey u forgot minus sign before 1 in second equation at the starting

OpenStudy (harsha19111999):

And i wrote wrong f(0)=1, f(1)=3

OpenStudy (harsha19111999):

f(1)-f(0) = 3-1 =2

OpenStudy (anonymous):

Oh yea I did so then would the answer be 2 because 3-1=2

OpenStudy (harsha19111999):

Yes you are correct

OpenStudy (anonymous):

Thank you I got it right:)

OpenStudy (harsha19111999):

You're welcome